The sum of the measures of the three angles of a triangle is 180 degrees .Find the three angles if two angles are the same and the third angle is 60 degrees more than the other two.
Answer by edjones(8007)

a=b
c=a+60
2a+a+60=180
3a=180-60
3a=120
a=40
b=40
c=100
.
Ed